# MAVERIX: Multimodal Audio-Visual Evaluation Reasoning IndeX

### v1

@@ Abstract
Frontier models have either been language-only or have primarily focused on vision and language modalities. Although recent advancements in models with vision and audio understanding capabilities have shown substantial progress, the field lacks a standardized evaluation framework for thoroughly assessing their cross-modality perception performance. We introduce MAVERIX (Multimodal Audio-Visual Evaluation Reasoning IndeX), a novel benchmark with 700 videos and 2,556 questions explicitly designed to evaluate multimodal models through tasks that necessitate close integration of video and audio information. MAVERIX uniquely provides models with audiovisual tasks, closely mimicking the multimodal perceptual experiences available to humans during inference and decision-making processes. To our knowledge, MAVERIX is the first benchmark aimed explicitly at assessing comprehensive audiovisual integration. Experiments with state-of-the-art models, including Gemini 1.5 Pro and o1, show performance approaching human levels (around 70% accuracy), while human experts reach near-ceiling performance (95.1%). With standardized evaluation protocols, a rigorously annotated pipeline, and a public toolkit (maverix-benchmark.github.io) [^1] Project page will be available soon., MAVERIX establishes a challenging testbed for advancing audiovisual multimodal intelligence.
Figure 1: MAVERIX contains over 2,500 audio-visual tasks spanning diverse domains with tight modality interdependence (left) and remains challenging, as state-of-the-art multimodal frontier models are consistently outperformed by average human participants (right).
Figure 2: Example questions and answers in the MAVERIX benchmark.
1 Introduction
@@
Human cognition seamlessly integrates visual and auditory information to reason, infer, and interact within dynamic environments. Replicating this ability in Multimodal Large Language Model (MLLM) systems remains a central challenge for AI, as autonomous agents must process complex audiovisual input to engage meaningfully with the world [32, 4, 1].
@@
Recent progress in multimodal foundation models has brought us closer to this goal, but current benchmarks fall short in assessing their abilities to reason with multimodal inputs. Most benchmarks focus on static images [11, 3], simple recognition, or tasks that can be solved through unimodal cues like transcripts [44, 13]. These benchmarks fail to probe the deeper, joint reasoning across modalities needed for real-world scenarios, such as interpreting social interactions or anticipating off-screen events [10, 36].
@@
A core obstacle in designing effective multimodal benchmarks is ensuring that tasks genuinely require multimodal understanding rather than allowing models to exploit unimodal shortcuts. Many existing video-language benchmarks reduce to visible-object recognition or dialogue parsing, sidestepping the need to synthesize audio-visual dependencies [40, 25, 29].
@@
To address this, we introduce MAVERIX, a benchmark designed to evaluate multimodal video-audio understanding through tasks that have tight modality interdependence. MAVERIX features four challenging tasks: situational awareness, contextual summarization, social sentiment analysis, and dynamic reasoning, covering 700 videos and over 2,500 carefully designed questions. These are constructed through a hybrid human-AI pipeline to ensure that solving them requires intertwined audiovisual synthesis, revealing the underlying multimodal reasoning capabilities of models applied to the benchmark tasks.
@@
Evaluations of state-of-the-art models, including Gemini 1.5 Pro [49], GPT-4o [39], and o1 [38], reveal significant gaps. Gemini 1.5 Pro, even with direct audio-video inputs, achieves 71.9% accuracy on multiple choice questions, marginally lower than human performance (80%). Open-ended responses further expose weaknesses in temporal reasoning and contextual understanding, with models averaging only 1.9/5 vs. human 2.79/5 in GPT-4o-judged scoring. Further, models that are not capable of processing raw audio and rely solely on transcripts perform even worse, highlighting the inadequacy of text-only proxies for rich audiovisual comprehension [37, 17].
@@
2 Benchmark Design and Constructions
Designing benchmarks for audiovisual understanding requires tasks that demand genuine cross-modal reasoning, not just unimodal pattern matching. In real-world scenarios, one modality often dominates, such as visuals in a silent event or audio in a podcast, but effective understanding depends on integrating cues across modalities. Existing benchmarks typically focus on narrow tasks or merge disjointed datasets, limiting their ability to evaluate general-purpose multimodal LLMs (MLLMs). Moreover, many video-language benchmarks rely on multiple-choice questions (MCQs) for their simplicity, but small answer spaces of the common four-option MCQs inflate scores through random guessing and fail to assess free-form reasoning — a core aspect of human-like comprehension.
Figure 3: Annotation workflow outlining the stages of data selection, review, and refinement.
To address these limitations, MAVERIX introduces a unified benchmark of diverse real-world tasks that systematically test the interdependence of modality while supporting a dual evaluation framework of eight-option MCQs and open-ended questions. The expanded eight-option format reduces guess bias (12.5% baseline), providing more reliable accuracy metrics, while open-ended tasks require models to generate nuanced, context-aware responses. For open-ended evaluation, we employ GPT-4o as an automated judge, scoring responses against human-curated references using a five-dimensional rubric: correctness of information, attention to detail, contextual understanding, temporal coherence, and consistency under paraphrasing. MAVERIX’s taxonomy and revealed performance gaps are shown in Fig. 1.
This hybrid design balances scalability and depth: MCQs provide standardized large-scale evaluation, while open-ended questions probe the models’ ability to synthesize and express complex audiovisual reasoning. Together, they create a robust assessment ecosystem tailored to intertwined multimodal understanding. We detail our pipeline used to generate MAVERIX’s 2,500+ high-quality questions and answers.
2.1 Dataset Generation Pipeline
Dataset Collection To ensure broad coverage of multimodal reasoning, we designed four domain hierarchies spanning agentic tasks, topics, video categories, and multimodal abilities, as shown in Fig. 2. Under agentic tasks, we defined seven key areas—social interactions, sentiment expression, information querying, sports analysis, gaming commentary, egocentric navigation, and multilingual contexts—each reflecting real-world scenarios that require complementary audiovisual cues. For example, gaming videos involve coordinating visual gameplay with voice chat strategies. In total, we annotated 155 fine-grained activity tags (e.g., “referee decision analysis” under sports) to capture diverse reasoning challenges (see Fig. 5).
Figure 4: Overview of the dataset distribution. The corresponding number of samples in each category is shown.
We sourced videos from YouTube-8M [2], MSR-VTT [53], UR-FUNNY-V2 [20], Ego4D [19], and AudioSet [18], balancing short (<1absent1<1< 1 min), medium (2–10 min), and long (10–65 min) formats. Each video includes subtitles, categorized audio (e.g., speech, environmental, artificial), and visual keyframes sampled at 1-second intervals. After deduplication and quality checks, MAVERIX comprises 700 videos with balanced coverage across domains and durations.
Multimodal Question-Answer Annotation To evaluate models’ ability to synthesize audiovisual information, we created 852 eight-option multiple-choice questions through a three-stage annotation protocol. A team of 11 researchers proficient in English (or another second language) with expertise in vision-language learning ensured high-quality, semantically precise questions.
The process involved (1) selecting videos and identifying clips requiring understanding multiple modalities, (2) drafting questions designed to be answerable only through joint audiovisual analysis, and (3) applying dual-validation checks to eliminate unimodal shortcuts by verifying that neither audio-only nor vision-only playback sufficed. The correct option is manually annotated, then paraphrased and modified into semantically different options with structurally similar alternative options. Surviving questions underwent peer review, with ambiguous cases resolved by consensus.
Figure 5: Dataset statistics. Left: distribution of video sources. Center: question and answer length across difficulty levels. Right: distribution of video durations for full-length and truncated clips.
For example, a valid question might ask, “Why did the mechanic abruptly stop speaking?” requiring both visual cues (e.g., discovering a leak) and audio cues (e.g., sudden silence). This protocol ensures MAVERIX’s QA pairs demand genuine modality interdependence, preventing reliance on any single modality.
@@ 2.1 Dataset Generation Pipeline
Quality Assurance To ensure the reliability of MAVERIX’s videos and annotations, we apply a two-stage quality control protocol combining expert review and model-based validation. Each QA pair undergoes four checks by a separate annotator, as illustrated in Fig. 3: (1) linguistic validity for clear and grammatical phrasing, (2) answerability of whether the question is resolvable via the video’s audiovisual content), (3) option integrity to ensure one correct answer with plausible distractors like semantically tangent or structurally identical options, and (4) modality interdependence, using cross-modal invalidation tests from Section 3.1 (e.g., disabling audio or video to detect shortcuts). For open-ended questions, reviewers also confirm that rephrased variants preserve meaning without overlapping with the ground-truth wording.
To guard against textual or visual biases, we test each question with GPT-4o for text-only and video keyframes + subtitles inputs. Questions answered correctly without full audiovisual reasoning are revised. For example, if GPT-4o answers “Why did the mechanic sigh?” from subtitles alone, we refine the question. Reviewers document model failures and work with annotators to iteratively resolve ambiguous cases. This combined human-model process ensures MAVERIX’s questions demand proper multimodal understanding.
2.2 MAVERIX Statistics
@@
MAVERIX comprises a diverse set of videos spanning 155 real-world scenarios across seven agentic task domains, such as social interactions, sentiment analysis, and egocentric understanding. These are complemented by hierarchies over topics (e.g., travel, technology), video categories (e.g., documentaries, vlogs), and multimodal abilities (e.g., temporal reasoning, emotion recognition), supporting broad and balanced evaluation. Each question is also tagged with a difficulty level reflecting its multimodal complexity as judged by human annotators. A detailed breakdown of these splits is provided in Fig. 4.
@@
The dataset contains 105.8 hours of video footage, with durations ranging from 10 seconds to 63 minutes. Videos are distributed across three duration categories: 16.8% short clips (<1 minute) for rapid context-switching tasks, 75.7% medium-length videos (1–10 minutes) for sustained narrative reasoning, and 7.4% long-form content (10–65 minutes) for testing temporal coherence. The average video length is 5.7 minutes, reflecting real-world use cases like troubleshooting tutorials or sports highlights.
@@
Each video is paired with 3–4 questions on average (2,556 total), including eight-option multiple-choice questions (MCQs) and open-ended prompts. Questions are evenly rebalanced across tasks to ensure a balanced evaluation. To mitigate positional bias, answer labels are uniformly redistributed across options (e.g., correct answers appear equally in positions A–H). Fig. 4 details the dataset’s domain composition, and Fig. 5 shows the duration distribution and question-type breakdown.
3 Experiments
3.1 Evaluation Protocol
@@
MAVERIX adopts a dual evaluation framework to assess multimodal LLMs (MLLMs) through eight-option multiple-choice questions (MCQs) and open-ended generation tasks.
@@
Evaluation is conducted under two settings: localized, where models access only the timestamped video segment relevant to each question, and global, where the full-length video is provided. The localized setting mitigates information leakage by limiting the context to the specific temporal window required for reasoning.
@@
For MCQs, we report both task-specific and overall accuracy, with answer choices uniformly distributed across positions (A–H) to reduce positional bias. Open-ended responses are evaluated using a GPT-4o-based scoring pipeline, adapted from Video-ChatGPT [35], which assesses outputs on five dimensions using a 0–5 scale. Scoring is calibrated on 200 human-annotated examples to ensure consistency, with penalties applied for failures in audiovisual reasoning, such as overlooking modality contradictions.
Results are aggregated across tasks (Tab. 1), with separate analyses for easy, medium, and hard videos to diagnose reasoning limitations. This protocol ensures reproducibility while addressing modality interdependence, which is the core challenge MAVERIX is designed to evaluate.
3.2 Baselines
@@
We evaluate MAVERIX on a diverse suite of 12 MLLMs, encompassing both commercial and open-source architectures, to assess their ability to reason over intertwined audiovisual modalities. Commercial models include Gemini 1.5 Pro [49], GPT-4o [39], o1 [38], Grok2-Vision [52], Claude Sonnet 3.5 [5], Nova-Lite [23], and Nova-Pro [23], while open source representatives feature InternVL2 [14], Qwen2.5-VL [6], LLaVA-OneVision [27], DeepSeek-VL2-Tiny [51], and DeepSeek-VL2-Small [51]. Gemini 1.5 Pro is uniquely equipped with native audiovisual processing, enabling direct ingestion of raw video-audio streams. For these models, we maximize temporal resolution by sampling frames at their maximum supported rates. However, most architectures, including GPT-4o and LLaVA-OneVision, require transcribed subtitles as text proxies for audio. To standardize inputs, we preprocess all videos using Whisper-base-v3 to extract time-synced subtitles, with the timestamps provided to the evaluated models.

### v2

# Liuyue Xie∗1  Avik Kuthiala∗1  George Z. Wei∗1 Ce Zheng1  Ananya Bal1  Mosam Dabhi1  Liting Wen1 Taru Rustagi1  Ethan Lai1  Sushil Khyalia1  Rohan Choudhury1

@@ Abstract
We introduce MAVERIX (Multimodal audiovisual Evaluation and Recognition IndeX), a unified benchmark to probe the video understanding in multimodal LLMs, encompassing video, audio, text inputs with human performance baselines. Although recent advancements in models with vision and audio understanding capabilities have shown substantial progress, the field lacks a standardized evaluation framework to thoroughly assess their cross-modality comprehension performance. MAVERIX curates 2,556 questions from 700 videos, in the form of both multiple-choice and open-ended formats, explicitly designed to evaluate multimodal models through questions that necessitate tight integration of video and audio information, spanning a broad spectrum of agentic scenarios. MAVERIX uniquely provides models with audiovisual questions, closely mimicking the multimodal perceptual experiences available to humans during inference and decision-making processes. To our knowledge, MAVERIX is the first benchmark aimed explicitly at assessing comprehensive audiovisual integration in such granularity. Experiments with state-of-the-art models, including Qwen 2.5 Omni and Gemini 2.5 Flash-Lite, show performance around 64% accuracy, while human experts reach near-ceiling performance of 92.8%, exposing a substantial gap to human-level comprehension. With standardized evaluation protocols, a rigorously annotated pipeline, and a public toolkit, MAVERIX establishes a challenging testbed for advancing audiovisual multimodal intelligence.
Introduction
@@
Human cognition seamlessly integrates visual and auditory information to reason, infer, and interact within dynamic environments. Replicating this ability in Multimodal Large Language Model (MLLM) systems remains a central challenge for AI, as autonomous agents must process complex audiovisual input to engage meaningfully with the world (Lin et al. 2023; Amirizaniani et al. 2024; tes 2024).
Figure 1: An illustration of our proposed benchmark, which includes highly audiovisual correlated questions and paraphrased questions, can be used to evaluate the model’s underlying comprehension abilities and their gaps to humans.
Figure 2: Example Agentic Categories and corresponding QAs in the MAVERIX benchmark.
@@
Recent progress in multimodal foundation models has brought us closer to this goal, but current benchmarks fall short in assessing their abilities to reason with multimodal inputs. Most benchmarks focus on static images (Chen et al. 2015; Agrawal et al. 2019; Li et al. 2025b), simple recognition, or questions that can be solved through unimodal cues, such as transcripts (Singh et al. 2019; Chen et al. 2024b). These benchmarks fail to probe the deeper, joint reasoning across modalities needed for real-world scenarios, such as interpreting social interactions or anticipating off-screen events (Chandrasegaran et al. 2024; Mangalam, Akshulakov, and Malik 2023).
@@
A core obstacle in designing effective multimodal benchmarks is ensuring that the questions genuinely require multimodal understanding rather than allowing models to exploit unimodal shortcuts or common sense from the training data. For benchmarks designed to expose the model understanding for highly multimodal data, their common adoption of a 4-way multiple-choice question for evaluation provides limited insights into the underlying interpretations (Li et al. 2025b; Hong et al. 2025). Many existing video-language benchmarks reduce to visible-object recognition or dialogue parsing, bypassing the need to synthesize audiovisual dependencies (Patraucean et al. 2023; Kesen et al. 2023; Li et al. 2024b).
@@
To address this, we introduce MAVERIX, a benchmark designed to evaluate multimodal video-audio understanding through questions that have tight modality interdependence. MAVERIX features questions from challenging agentic scenario categories: factual recall, causal understanding, sentiment analysis, temporal recall, situational awareness, context understanding, social interaction understanding, and emotional synthesis, covering 700 videos and 2,556 carefully designed questions. These are constructed through a hybrid human-AI pipeline to ensure that solving them requires intertwined audiovisual synthesis, revealing the underlying multimodal reasoning capabilities of models applied to the benchmark questions.
@@
Evaluations of state-of-the-art proprietary and open-source models of different sizes, including Gemini 2.0/2.5 Flash-Lite(FL)(Team and Petko Georgiev 2024), GPT-4o (OpenAI and Aaron Hurst 2024), and Qwen 2.5 Omni (Xu et al. 2025), reveal significant gaps as shown in Fig. 1. Gemini 2.5 Flash-Lite, even with direct audio-video inputs, achieves 54.7% accuracy on multiple choice questions, significantly lower than human performance (92.8%). Open-ended responses further expose weaknesses in temporal reasoning and contextual understanding, with models averaging only 1.9/5 vs. human 2.79/5 in GPT-4o-judged scoring. Further, models that are not capable of processing raw audio and rely solely on transcripts perform even worse, highlighting the inadequacy of text-only proxies for rich audiovisual comprehension (Ning et al. 2023; Fu et al. 2024).
@@
Benchmark Design and Construction
MAVERIX challenges multimodal large language models (MLLMs) to integrate audio and visual evidence under realistic conditions. This section elaborates on four key aspects: (i) the motivation behind our design, (ii) the dataset construction pipeline, (iii) the dual-format evaluation protocol, and (iv) dataset statistics. Fig. 3 visualizes the pipeline; Tab. 1 compares the benchmark components with the relevant works; Tab. 2 summarizes key dataset statistics.
Benchmark #Vid. Med. Len. (s) #Q Mod. MCQ. #Div. Diff. Shortcut Human OE. MSRVTT-QA 2,990 15 72,821 V 4-MCQ 1 MSVD-QA 504 9 50,505 V 4-MCQ 1 ActivityNet-QA 5,800 15 58,000 V 4-MCQ 1 How2QA 1,517 11 71,812 V+S+A 4-MCQ 1 AutoEval-Video 327 32 450 V 4-MCQ 1 TempCompass 410 11 1,540 V 4-MCQ 2 Video-MME 900 1,072 2,700 V+S 4-MCQ 3 OmniBench 1142 I+A 4-MCQ 3 WorldSense 1,662 141.1 3,172 V+S+A 4-MCQ 3 HourVideo 500 2,742 12,976 V+S 4-MCQ 2 MAVERIX 700 106 2,556 V+S+A 8-MCQ 7 MAVERIX-Long 700 345 2,556 V+S+A 8-MCQ 7
Table 1: Comparison with prior video-question benchmarks. Mod.: V (video), S (subtitles), A (audio). Ans.: 4-MCQ, “8-MCQ+OE” (eight-option plus open-ended). Diff.: crowdsourced or expert difficulty labels. #Div.: Number of division types. Shortcut: dataset validated against audio-only / video-only ablations. Human: ✓ if a benchmark reports any human baseline.
Design Motivation and Principle
While previous video-understanding benchmarks curate multiple-choice question-answer pairs over different topics, some still suffer unimodal shortcuts and are limited in exposing models’ underlying biases. For evaluating the models’ multimodal understanding abilities, we source videos that capture a wide range of temporal events, spatial motions, and audiovisual correlations. While sourcing the videos and constructing the multiple choice and open-ended question-answer pairs, we follow the following design principles.
Avoid Unimodal Shortcut. Many existing image and video based question answering benchmarks (e.g., TVQA, MSR‑VTT Q&A) contain questions that can be answered from captions or a single salient frame, enabling unimodal shortcuts.
Wide range of evaluation dimensions. Deploying MLLMs in the real world requires that the models understand and handle scenarios with different skills. Single‑skill benchmarks do not reflect the breadth of reasoning required in open‑world settings. The curated benchmark measures the models in six evaluation dimensions, covering acoustic understanding, agentic skills, understanding by broad and sub-topic taxonomy, temporal understanding, and multimodal synthesis abilities.
Prevention of guess inflation. We design a hybrid of eight-way multiple-choice, and open-ended QA for the benchmark, such that by-passing the questions with model-inherent biases can be evidently exposed. The hybrid design evaluates the models’ actual abilities to interpret the input sources in different modality settings and provides a fair evaluation of their capabilities.
Dataset Generation Pipeline
Video Collection. We primarily source our video content from five datasets: YouTube‑8M(Abu-El-Haija et al. 2016), MSR‑VTT(Xu et al. 2016), UR‑FUNNY‑V2(Hasan et al. 2019), Ego4D(Grauman et al. 2021), and AudioSet(Gemmeke et al. 2017). YouTube‑8M(Abu-El-Haija et al. 2016) and AudioSet(Gemmeke et al. 2017) are large-scale datasets covering a wide range of taxonomies, where the videos have strong audiovisual correlations. MSR-VTT comprises of high quality video descriptions designed for video translation QAs. UR-FUNNY-V2 exhibits videos of different emotional states, challenging the models in their sentimental understanding. Lastly, Ego4D consists of egocentric, long-duration videos to probe model’s understanding to daily interactions for agentic scenarios. The videos are selected by human annotators according to the principles described above, ensuring a thorough distribution across the topics, with different durations: short with <1<1 min, medium with 1-10 mins, and long with 10-65 mins. Each video is processed and accompanied by subtitles generated with Whisper-v3 (Radford et al. 2022) to ensure a fair evaluation on video-text models without audio-support.
Initial Question Answering Annotation. A team of 8 expert annotators engaged in the initial question-answering pair curation. The annotators provided at least one question answer pair to each video to generate the initial ground truths. Then the same pair is expanded into eight-way multiple choice question with alternative distractive answers.
Shortcut Removal and Validation. Following the initial annotation, we use a semi-automated approach to validate difficulties with MLLMs and refine the questions to avoid any potential shortcuts. Each question undergoes three ablation tests with GPT‑4o-mini and Gemini 2.0-FL: text‑only, video‑only, and videos+subtitles. If any ablation yields the correct answer for both models, the item is flagged and revised to reduce reliance on unimodal cues. All revisions are logged, and the final set is approved after a second expert pass. For example, a valid question might ask, “Why did the mechanic abruptly stop speaking?” requiring both visual cues (e.g., discovering a leak) and audio cues (e.g., sudden silence). This protocol ensures MAVERIX’s QA pairs demand genuine modality interdependence, preventing reliance on any single modality.
The difficulty labels were crowd-sourced through Amazon MTurk service (mtu 2005) with 219 unique participants for gauging common consensus, and are determined based on the subtlety of cross-modal cues, the depth of understanding required, and the ease of locating relevant information in the video. The human performance evaluations were gathered through MTurk with 382 participants answering a 1/3 subset of the MCQs and open-ended questions.
Statistic Audio Type Agentic Categories Topic Domain Overall
Sub-class of QA
Mixed
Sound
Speech
Artificial
Sound
Information
Querying
Egocentric
Agent
Sentiment
Analysis
Humanity &
Society
Business &
Commerce
Agentic Abilities
Causal Relationship 66 111 15 75 27 33 27 21 201
Emotional Inference 51 57 18 66 21 27 27 12 129
Factual Recall 516 690 60 672 171 186 210 120 1311
Situational Understanding 27 27 10 33 9 12 12 6 70
Context Understanding 309 414 45 237 147 138 111 84 771
QA Lengths
Question 11.26 11.63 9.06 10.31 10.41 10.00 10.09 10.51 11.28
Options 11.76 10.52 11.99 9.92 12.29 12.89 10.30 12.54 11.13
Open-ended answer 13.30 12.69 11.68 10.46 12.77 13.51 12.73 12.12 12.85
Subtitle length 440.79 682.24 419.87 485.08 1257.19 360.40 351.32 488.34 558.06
Video–Audio Statistics
Media Length 319.98 381.24 327.25 315.77 1039.18 289.09 222.42 259.69 352.63
Media min. Length 6.15 10.03 5.57 5.57 10.04 5.57 5.57 10.15 5.57
Media max. Length 6620.63 4427.76 3205.50 6620.63 6620.63 3851.93 1800.17 511.23 6620.63
Media std. 527.50 681.63 482.65 603.49 1341.37 439.54 268.20 117.12 610.82
Table 2: Statistics from the included data. Agentic Categories: counts per category. QA Lengths: mean word counts for questions, mean per-option length (MCQs), and open-ended answers (computed only when options are absent). Video–Audio Statistics: duration in seconds (mean, min, max, std).
Quality Assurance
@@
To ensure the reliability of MAVERIX’s videos and annotations, each QA pair undergoes four checks by an expert annotator, as illustrated in Fig. 3: (1) linguistic validity for clear and grammatical phrasing, (2) answerability of whether the question is resolvable via the video’s audiovisual content), (3) option integrity to ensure one correct answer with plausible distractors like semantically tangent or structurally identical options, and (4) modality interdependence, using cross-modal invalidation tests from Section Evaluation Protocol (e.g., disabling audio or video to detect shortcuts). For open-ended questions, reviewers also confirm that rephrased variants preserve meaning without overlapping with the ground-truth wording.
Figure 3: The framework to construct annotation sets with hybrid annotator and MLLM-as-judge quality assurance.
Dual‑Format Evaluation
[…]
Open‑Ended Generation. Every clip is paired with at least one semantically unique free-form prompt, ranging from causal explanation to future prediction, requiring natural‑language output. For each unique question, we paraphrase the question and prompt the model again to test its robustness against paraphrasing. The open-ended responses are graded by GPT‑4o on a five‑factor rubric covering factual correctness, attention to detail, contextual grounding, temporal coherence, and paraphrase robustness.
Together, the MCQ and generation tracks marry scalability, through high-throughput accuracy metrics, with depth, by exposing weaknesses in explanation quality.
MAVERIX Statistics
@@
MAVERIX comprises a diverse set of videos spanning 155 real-world scenarios across evaluation dimensions including agentic categories, topic domains, sub-topic domains, audio categories, duration, and difficulty. These are complemented by hierarchies over topics (e.g., travel, technology), video categories (e.g., documentaries, vlogs), and multimodal abilities (e.g., temporal reasoning, emotion recognition), supporting broad and balanced evaluation. Each question is also tagged with a difficulty level judged by human annotators. A detailed breakdown of these splits is provided in Tab.  2.
@@
The dataset contains 105.8 hours of video footage, with durations ranging from 10 seconds to 63 minutes. Videos are distributed across three duration categories: 16.8% short clips (<1 minute) for rapid context-switching understanding, 75.7% medium-length videos (1-10 minutes) for sustained understanding, and 7.4% long-form content (10-65 minutes) for testing temporal coherence. The average video length is 352.63 seconds, with an std. of 610.82, suggesting that the included medias have a diverse distribution in lengths. Constructing a audiovisual benchmark with diverse media lengths reflects real-world use cases for MLLMs and sufficiently challenges their ability to generalize.
@@
Each video is paired with 3 to 4 questions on average with 2,556 in total, including 852 eight-option multiple-choice questions (MCQs) and 1704 open-ended prompts, with examples shown in Fig. 2. Questions span over the evaluation dimensions to ensure a thorough evaluation. To mitigate positional bias, answer labels are uniformly redistributed across options.
Experiments
Evaluation Protocol
@@
MAVERIX adopts a dual evaluation framework to assess multimodal LLMs (MLLMs) through eight-option multiple-choice questions (MCQs) and open-ended response generation.
@@
Evaluation is conducted under two settings: localized, where models access only the timestamped video segment relevant to each question, and global (MAVERIX-Long), where the full-length video is provided. The localized setting limits the context to the specific temporal window required for understanding, whereas providing the full videos demands the models to localize the required information from the haystack of frames.
@@
For MCQs, we report both split-specific and overall accuracy, with answer choices uniformly distributed across positions (A-H) to reduce positional bias. Open-ended responses are evaluated using a GPT-4o scoring pipeline, adapted from Video ChatGPT (Maaz et al. 2024), which assesses the output in five dimensions on a scale of 0-5. The results are aggregated across modalities (Tab. 3), with separate analyses for easy, medium, and hard videos to diagnose comprehension limitations in Tab. 4. We also report the model’s open-ended response qualities and token counts in Tab. 6. Our proposed evaluation protocol ensures reproducibility while addressing modality interdependence and human performance baselines.
Baselines
@@
We evaluate MAVERIX on a diverse suite of 17 MLLMs, encompassing both proprietary and open-source models, to assess their ability to reason over intertwined audiovisual modalities. Proprietary models include Gemini 2.0-FL (Team and Petko Georgiev 2025), Gemini 2.5-FL (Team and Petko Georgiev 2025), GPT-4o (OpenAI and Aaron Hurst 2024), Grok4 (xAI 2025), Claude Sonnet 3.5 (Anthropic 2024), Nova-Lite (Intelligence 2024), and Nova-Pro (Intelligence 2024). While open source representatives feature Ola (Liu et al. 2025b), EgoGPT (Yang et al. 2025), VITA 1.5 (Fu et al. 2025), Qwen 2.5 Omni (Xu et al. 2025), InternVL2 (Chen et al. 2024c), Qwen2.5-VL (Bai et al. 2025), LLaVA-OneVision (Li et al. 2024a), DeepSeek-VL2-Small (Wu et al. 2024). Among them, Ola, EgoGPT, VITA 1.5, Qwen 2.5 Omni and Gemini are equipped with native audiovisual processing, enabling direct ingestion of raw video-audio streams. For the tested models, we maximize temporal resolution by sampling frames at their maximum supported rates. However, most architectures, including GPT-4o and LLaVA-OneVision, require transcribed subtitles as text proxies for audio. To standardize inputs, we preprocess all videos using Whisper-v3(Radford et al. 2022) to extract time-synced subtitles, with the timestamps provided to the evaluated models.
[…]
Model
Audio
Support
Size Arch Recipe Unimodal Analysis Multimodal Performance
[…]
Human $44.3$ $41.7$ 81.4 81.4 86.4 92.8 92.8 +11.4+11.4
EgoGPT-7B 7B Dual-Tower SFT $29.9$ $43.0$ $45.2$ $45.2$ $55.0$ $45.2$ $55.0$ +9.8+9.8
Ola-7B 7B Tri-Tower SFT 49.4 $43.9$ $37.6$ $49.5$ $49.6$ $53.1$ $53.1$ +3.6+3.6
VITA 1.5 7B Dual-Tower SFT $32.4$ $43.4$ $20.2$ $43.5$ $22.3$ $18.5$ $22.3$ 21.2-21.2
Qwen 2.5 Omni 7B Dual-Tower SFT+RL $46.5$ $41.4$ $35.4$ $46.5$ $57.9$ $49.5$ $57.9$ +11.4+11.4
Qwen-2-VL 7B ViT-LLM SFT $43.0$ $48.0$ $48.0$ $57.5$ $57.5$ +9.5+9.5
Qwen-2.5-VL 7B ViT-LLM SFT $40.3$ $46.9$ $46.9$ $55.3$ $55.3$ +8.4+8.4
InternVL2 8B ViT-LLM SFT $24.1$ $26.3$ $26.3$ $33.1$ $33.1$ +6.8+6.8
LLaVA-OneVision 7B SigLIP-LLM SFT $44.5$ $46.8$ $46.8$ $55.6$ $55.6$ +8.8+8.8
DeepSeekVL2-small 2.8B Hybrid Enc.-MoE SFT $34.3$ $33.2$ $34.3$ $42.4$ $42.4$ +8.1+8.1
Gemini 2.0-FL $43.8$ $38.0$ $42.1$ $43.8$ $41.1$ $50.2$ $50.2$ +6.4+6.4
Gemini 2.5-FL $44.8$ $47.7$ $48.8$ $48.8$ $56.7$ $54.7$ $56.7$ +7.9+7.9
Claude Sonnet 3.5 $55.0$ $42.0$ $55.0$ 64.1 64.1 +9.1+9.1
GPT-4o 55.3 $54.3$ 55.3 $64.0$ $64.0$ +8.7+8.7
Grok 4 $41.8$ $54.5$ $54.5$ $59.4$ $59.4$ +4.9+4.9
GPT-4o-mini $45.4$ $35.5$ $45.4$ $50.0$ $50.0$ +4.6+4.6
NOVA-Lite $40.4$ $39.7$ $40.4$ $51.0$ $51.0$ +10.6+10.6
NOVA-Pro $46.6$ $45.4$ $46.6$ $55.8$ $55.8$ +9.2+9.2